OpenStudy (anonymous):

What is the slope of the line passing through the points (-2,4) and (3,6)

OpenStudy (jdoe0001):

http://s3.hubimg.com/u/1923390_f496.jpg

OpenStudy (texaschic101):

slope (m) = (y2 - y1) / (x2-x1).....set 1 (-2,4)..set 2 (3,6) m =(6 - 4)/(3 -(-2) m =(6 - 4)/(3 + 2) can you take it from here ?